OpenStudy (anonymous):

Find the x- and y-intercepts for the equation. Then graph it. 5x - 20y = 20

OpenStudy (anonymous):

A. (0, 4), (-1, 0)

OpenStudy (anonymous):

B. (0, -1), (4, 0)

OpenStudy (anonymous):

C. (0, -4), (1, 0)

OpenStudy (anonymous):

5x - 20y = 20 -20y=-5x +20 y = 1/4x -1

OpenStudy (anonymous):

plug in and you will know

OpenStudy (anonymous):

D. (0, -1), (-4, 0)

OpenStudy (anonymous):

try the numbers you put in the equation to get the answer example A) (0, 4), (-1, 0) y = 1/4x -1 y = 1/4(0) -1 ≠ 4 so its not A

OpenStudy (anonymous):

any luck?

OpenStudy (anonymous):

no im lost

OpenStudy (anonymous):

is c?

OpenStudy (anonymous):

nope (0, -4), (1, 0) y = 1/4x -1 y = 1/4(-4) -1 = -1 - 1= -2 ≠ 0 so its not C either

OpenStudy (anonymous):

Then its D